# Zero-Shot Strategies for Length-Controllable Summarization

### v1

@@ 1 Introduction
Studies on length controllability have generally been narrow, focusing on a single control scenario defined by a single length measure, such as word count (Juseon-Do et al., 2024; Yuan et al., 2024). Moreover, most approaches rely on fine-tuning models with additional training data or modifying model architectures–like length-aware attention mechanisms (Urlana et al., 2024). Although these methods can be effective, their dependence on model internals and substantial computational resources limits their practicality in real-world settings where only inference APIs are accessible.
@@
  • We explore combinations of proposed methods yielding highly effective ’recipes’ for length-controllable summarization with LLMs.
@@ 3.3.1 Word Target Adjustment
For word targets, the empirically determined coefficients (see Appendix F.3 for details) are:
@@ 4 Experimental Setup
We conducted a systematic evaluation of both the inherent length control capabilities of the Llama-3-8B-Instruct model (Dubey et al., 2024) and potential improvements through our proposed methods, using the YTSeg dataset (Retkowski and Waibel, 2024). More than 5M summaries were generated during the evaluation, and to support further analysis, we release all generated artifacts [^1] https://github.com/scientiaetveritas/llama3-length-controllable-summarization-evaluation.
@@
The YTSeg dataset, comprising $19,299$ video transcripts, was selected for its diversity in length, structure, and domain. Transcript lengths in this dataset range from a few hundred to over ten thousand words. This variability ensured a robust assessment of length controllability across heterogeneous inputs.
@@
Five length measures were examined with various targets, as summarized in Table 2. This comprehensive set of targets enables a multifaceted assessment of length control across varying granularities.
@@
Table 3: Influence of Response Priming (RP) and Sampling Temperature (Temp.) on Length Adherence
@@ 4.1 Experimental Conditions
    Response Priming: Restating the instruction by prefilling the response to improve instruction-following (Figure A3).
@@
    Revision Prompt: An automated follow-up instruction to refine summaries that do not meet length constraints, with additional response priming (Figure A3)
@@
To evaluate the performance of our length control methods, we employ six metrics: Exact Match (EM), Length Compliance (LC), Length Deviation (LD), Compression Rate (CR), Perplexity (PPL), and AlpacaEval. EM quantifies the proportion of generated lengths that precisely match their targets, while LC measures adherence within a specified tolerance. LD quantifies the absolute difference between observed and target lengths, and CR provides insight into data compression by comparing target and observed lengths. PPL helps assess model uncertainty, providing insight into generation quality. In line with recent work on length-controllable summarization Yuan et al. (2024), we also use the AlpacaEval protocol Dubois et al. (2023) as an LLM-as-a-Judge metric to assess overall generation quality. For detailed definitions of these metrics, please refer to Appendix C.
Revision Trigger.
Response Priming and Temperature.
The impact of response priming (RP) and sampling temperature $\tau\in\{0.3,0.7,1.0\}$ on length adherence was initially investigated. For subsequent experiments, we assumed the usage of response priming and a temperature of 0.7, based on their performance in preliminary tests.
@@
Figure 2: Analysis of length deviation across different parameters with data downsampled for visualization. Target summary lengths ranged from 25 to 200 words. Smoothed trend line using local regression (LOESS).
Context Length.
@@
Table 4: Empirical Analysis of Length Controllability Across Measures
@@ 5 Results
Influence of RP and Temperature.
We first investigate how response priming and sampling temperature affect the model’s adherence to length constraints. As shown in Table 3, incorporating RP significantly improves instruction-following and improves both LC and LD across different temperatures. For a target length of 50 words, LC improves from 75.6% without RP to 79.7% with RP at a temperature of 0.7. Lower sampling temperatures slightly enhance length control consistency, but the combination of RP and a temperature of 0.7 provides a good balance between adherence and output diversity. Therefore, we use RP and a temperature of 0.7 in subsequent experiments.
@@
Analysis of qualitative length quantifiers reveals the model’s attempt to adjust summary lengths accordingly (Table 5). These quantifiers yield a range of compression rates and output lengths, with “short” summaries averaging $127.2\pm 44.9$ words ($7.0\pm 4.8$% CR) and “long” summaries reaching $343.0\pm 83.9$ words ($19.8\pm 13.5$% CR). Notably, heteroscedasticity in output metrics across quantifiers indicates inconsistent summarization control. Figure 3 further illustrates non-linear trends and variability in both summary length and compression rate relative to input length. This inconsistency, combined with heteroscedastic results, suggests that qualitative length quantifiers provide imprecise control, limiting their effectiveness for length-controllable summarization tasks.
@@
Table 5: Influence of Quantifiers on Summary Length
@@
Additional results in Appendix J.3 provide insight into length controllability and biases across various models. Our analysis compares different LLaMA versions and other LLMs in zero-shot scenarios, highlighting that observed patterns of under- and over-generation and preference for certain measures are broadly consistent across architectures. This suggests that the length control challenges and model behaviors identified in our study are not exclusive to a single model configuration, reinforcing the relevance of our findings for a broad range of LLMs.
@@
Table 6: Performance of different methods across character, token, and word length measures
@@ 5.1 Effectiveness of Proposed Methods
Figure 4: Length Compliance with the Number of Revisions or Samples
@@
Table 7: Win Rates (%; ↑) against the baseline for different measures and target lengths. Results obtained using GPT-4o as judge, evaluating 100 pairs per condition.
@@
Our methods preserve—and in some cases enhance—summary quality. Since they do not modify model parameters, significantly alter prompts, or impose restrictive decoding constraints, we expected minimal impact on summary quality. This expectation is supported by our perplexity analysis in Table A1, which shows no degradation, suggesting that key content quality features such as coherence and fluency are preserved. The only notable difference is a lower perplexity in the LA approach, implying that summaries generated via word measure may achieve higher quality than those based on finer-grained measures like tokens or characters. This finding is further validated through AlpacaEval (see Table 7), where GPT-4o as a judge ranks our controlled summaries comparable to baseline outputs (SF: 50.3%, AR: 50.8%, TA: 53.5%), with LA again achieving a notably higher win rate of 61.9%. These results confirm that our interventions maintain or enhance content quality while substantially improving length control.
@@ Appendix A Prompt Templates
Figure A2: Prompting with Response Priming
@@ Appendix D Perplexity Scores
Appendix E AlpacaEval Evaluation Details
@@ Appendix I Results on Sampled Revisions
Appendix J Cross-Model Comparison
@@
Figure A4: Analysis of length deviation across different parameters and models with data downsampled for visualization. Target summary lengths ranged from 25 to 200 words. Smoothed trend line using local regression.
@@
Non-LLaMA models (Mistral-Nemo-Instruct-2407, gemma-2-9b-it, and aya-expanse-8b) tend to overgenerate content for shorter targets. For example, at 50-word targets, their compliance can be as low as 14.5%, with some summaries substantially exceeding the requested length. However, as target lengths increase beyond approximately 125 words, these models—like LLaMA variants—begin to undergenerate, often undershooting the requested length by tens of words. This pattern suggests a fundamental and broad-based difficulty for current language models: they struggle to maintain precise length control as output requirements become more substantial.

### v2

@@ 1 Introduction
Studies on length controllability have generally been narrow, focusing on a single control scenario defined by a single length measure, such as word count (Juseon-Do et al., 2024; Yuan et al., 2024). Moreover, most approaches rely on fine-tuning models with additional training data or modifying model architectures—like length-aware attention mechanisms (Urlana et al., 2024). Although these methods can be effective, their dependence on model internals and substantial computational resources limits their practicality in real-world settings where only inference APIs are accessible.
@@
  • We explore combinations of proposed methods yielding highly effective ’recipes’ for length-controllable summarization (LCS) with LLMs.
@@ 3.3.1 Word Target Adjustment
For word targets, the empirically determined coefficients (see Appendix E for details) are:
@@ 4 Experimental Setup
We conducted a systematic evaluation of both the inherent length control capabilities of the Llama-3-8B-Instruct model (Dubey et al., 2024) and potential improvements through our proposed methods, using the YTSeg (Retkowski and Waibel, 2024) and CNN/DM See et al. (2017) datasets. More than 6M summaries were generated during the evaluation, and to support further analysis, we release all generated artifacts [^1] https://huggingface.co/datasets/retkowski/length-controllability-evaluation.
@@
The YTSeg dataset ($19,299$ video transcripts) was selected for its diversity in length, structure, and domain. Transcript lengths in this dataset range from a few hundred to over ten thousand words. This variability ensured a robust assessment of length controllability across heterogeneous inputs. Additionally, we evaluate on the CNN/DM test set (11,490 articles) as a well-established benchmark in summarization.
@@
Five length measures were examined with predefined targets, as shown in Table 2, and reference summary lengths from CNN/DM. This comprehensive set of targets enables a multifaceted assessment of length control across varying granularities.
@@
Table 3: Influence of Prompt Prefilling (PP) and Sampling Temperature (Temp.) on Length Adherence on YTSeg
@@ 4.1 Experimental Conditions
    Prompt Prefilling: Restating the instruction by prefilling the response to improve instruction-following (Figure A3).
@@
    Revision Prompt: An automated follow-up instruction to refine summaries that do not meet length constraints, with additional prompt prefilling (Figure A3)
@@
To evaluate the performance of our length control methods, we employ seven metrics: Exact Match (EM), Length Compliance (LC), Length Deviation (LD), Compression Rate (CR), Perplexity (PPL), AlpacaEval, and ROUGE. EM quantifies the proportion of generated lengths that precisely match their targets, while LC measures adherence within a specified tolerance. LD quantifies the absolute difference between observed and target lengths, and CR measures data compression by comparing target and observed lengths. PPL helps assess model uncertainty, providing insight into generation quality. We employ ROUGE as a classic measure of summary quality. In line with recent work on LCS Yuan et al. (2024), we also use the AlpacaEval protocol Dubois et al. (2023) as an LLM-as-a-Judge metric to assess overall generation quality. For detailed definitions of these metrics, please refer to Appendix C.
@@
Figure 2: Analysis of length deviation across different parameters on YTSeg with data downsampled for visualization. Target summary lengths range from 25 to 200 words. Smoothed trend line using local regression (LOESS).
@@
Table 4: Empirical Analysis of Length Controllability Across Measures on YTSeg
@@ 5 Results
Influence of PP and Temperature.
We first investigate how prompt prefilling and sampling temperature affect the model’s adherence to length constraints. As shown in Table 3, incorporating PP significantly improves instruction-following and improves both LC and LD across different temperatures. For a target length of 50 words, LC improves from 75.6% without PP to 79.7% with PP at a temperature of 0.7. Lower sampling temperatures slightly enhance length control consistency, but the combination of PP and a temperature of 0.7 provides a good balance between adherence and output diversity. Therefore, we use PP and a temperature of 0.7 in subsequent experiments.
@@
Analysis of qualitative length quantifiers reveals the model’s attempt to adjust summary lengths accordingly (Table 5). These quantifiers yield a range of compression rates and output lengths, with “short” summaries averaging $127.2\pm 44.9$ words ($7.0\pm 4.8$% CR) and “long” summaries reaching $343.0\pm 83.9$ words ($19.8\pm 13.5$% CR). Notably, heteroscedasticity in output metrics across quantifiers indicates inconsistent control. Figure 3 further illustrates non-linear trends and variability in summary length and CR relative to input length. This inconsistency, combined with heteroscedastic results, suggests that qualitative quantifiers provide imprecise control, limiting their effectiveness for LCS tasks.
@@
Table 5: Influence of Quantifiers on Summary Length on YTSeg
@@
Additional results in Appendix J provide insight into length controllability and biases across various models. Our analysis compares different LLaMA versions and other LLMs in zero-shot scenarios, highlighting that observed patterns of under- and over-generation and preference for certain measures are broadly consistent across architectures. This suggests that the length control challenges and model behaviors identified in our study are not exclusive to a single model configuration, reinforcing the relevance of our findings for a broad range of LLMs.
@@
Table 6: Performance of different methods across character, token, and word length measures on YTSeg
@@ 5.1 Effectiveness of Proposed Methods
Figure 4: Length Compliance with the Number of Revisions or Samples on YTSeg
@@
Characters Tokens Words
Method LC LD R1 R2 RL LC LD R1 R2 RL LC LD R1 R2 RL
Baseline 17.7 $93.6$ $28.6$ $9.0$ $24.6$ 3.3 $60.3$ $30.8$ $10.2$ $25.1$ 67.7 $4.4$ $30.4$ $9.4$ $25.5$
LA 58.3 $31.2$ $30.5$ $9.5$ $25.7$ 61.9 $6.3$ $31.1$ $9.8$ $26.1$
AR 47.1 $50.4$ $29.2$ $8.6$ $24.3$ 18.4 $21.2$ $31.3$ $9.8$ $25.7$ 92.7 $2.8$ $30.2$ $9.2$ $25.3$
SF 39.2 $62.9$ $29.0$ $8.8$ $24.6$ 9.1 $39.0$ $31.3$ $10.1$ $25.6$ 94.0 $2.0$ $30.5$ $9.3$ $25.5$
SF-AR 64.4 $37.0$ $29.5$ $8.9$ $24.7$ 29.8 $16.2$ $31.5$ $10.0$ $26.0$ 98.9 $1.7$ $30.5$ $9.3$ $25.5$
SR 79.4 $23.4$ $29.4$ $8.7$ $24.5$ 46.2 $10.7$ $31.2$ $9.8$ $25.8$ 99.7 $1.6$ $30.5$ $9.3$ $25.5$
LA-AR 71.7 $33.9$ $30.0$ $9.0$ $25.0$ 70.3 $14.0$ $30.6$ $9.4$ $25.5$
LA-SF 89.0 $15.3$ $30.1$ $9.1$ $25.2$ 89.5 $3.1$ $30.7$ $9.5$ $25.6$
LA-SF-AR 93.1 $16.0$ $30.1$ $9.1$ $25.2$ 92.4 $4.6$ $30.7$ $9.5$ $25.7$
LA-SR 95.3 $13.7$ $30.1$ $9.1$ $25.2$ 94.2 $3.8$ $30.6$ $9.5$ $25.6$
Table 7: Performance of the proposed methods on the CNN/DM dataset, with reference length as target length
@@
Table 8: Win Rates (%; ↑) against the baseline for different measures and target lengths on YTSeg. Evaluating 100 pairs per condition, using GPT-4o as judge.
@@
Our methods preserve—and in some cases enhance—summary quality. Since they do not modify model parameters, significantly alter prompts, or impose restrictive decoding constraints, we expected minimal impact on quality. This expectation is supported by multiple lines of evidence. First, our perplexity analysis in Table A1 shows no degradation, suggesting that key content quality features such as coherence and fluency are preserved. The only notable difference is a lower perplexity in the LA approach, implying that summaries generated via word measure may achieve higher quality than those based on finer-grained measures like tokens or characters. This finding is further validated through AlpacaEval (see Table 8), where GPT-4o as a judge ranks our controlled summaries comparable to baseline outputs (SF: 50.3%, AR: 50.8%, TA: 53.5%), with LA again achieving a notably higher win rate of 61.9%. Lastly, the ROUGE scores in Table 7 largely remain stable across all methods, with some approaches slightly improving the performance and achieving substantially better length control. These results confirm that our interventions maintain or enhance content quality while providing superior length control.
@@ Appendix A Prompt Templates
Figure A2: Prompting with Prompt Prefilling
@@ Appendix D Perplexity Scores
Appendix E Additional Experimental Details
Revision Trigger.
Prompt Prefilling and Temperature.
The impact of prompt prefilling and sampling temperature $\tau\in\{0.3,0.7,1.0\}$ on length adherence was initially investigated. For subsequent experiments, we assumed the usage of PP and a temperature of 0.7, based on their performance in preliminary tests.
Context Length.
AlpacaEval Evaluation.
@@ Appendix I Results on Sampled Revisions
Appendix J Cross-Model Experiments
@@
Figure A4: Analysis of length deviation across different parameters and models with data downsampled for visualization. Target summary lengths range from 25 to 200 words. Smoothed trend line using local regression.
@@
Non-LLaMA models (Mistral-Nemo-Instruct-2407, gemma-2-9b-it, and aya-expanse-8b) tend to overgenerate content for shorter targets. For example, at 50-word targets, their compliance can be as low as 14.5%, with some summaries substantially exceeding the requested length. However, as target lengths increase beyond approximately 125 words, these models—like LLaMA variants—begin to undergenerate, often undershooting the requested length by tens of words. This pattern suggests a fundamental and broad-based difficulty for current language models: they struggle to maintain precise length control as target length increases.